# Generalization in Reinforcement Learning for Radio Access Networks

### v2

@@ 5.1 Scaling Training in Network Simulators
In this work, we implement a large-scale distributed RL algorithm, illustrated in Figure 2. The algorithm is based on the distributed prioritized experience replay architecture proposed in Horgan et al. (2018), which decouples inference from learning, enables parallel inference, and significantly improves data collection capabilities. The architecture comprises three key components: a single learner operating on a graphics processing unit (GPU); multiple independent actors running concurrently on central processing units; and a centralized, yet sharded, prioritized replay memory. A sharded replay memory partitions the experience replay buffer into several smaller segments, or shards, each storing and managing a portion of the collected experiences independently—often distributed across multiple threads, processes, or machines—to enhance scalability and throughput. Each actor concurrently interacts with a group of simulations hosted by different CPUs, further accelerating the data collection process. Furthermore, each actor employs a distinct exploration policy to enrich the diversity of data. Experiences generated by each actor through interactions with its environment instances are asynchronously contributed to the replay memory shard.
@@ 5.2 Working Principles
The interaction among actors, shards, and the learner—illustrated in Figure 2—proceeds as follows:
@@
The learner periodically queries each shard to retrieve its total priority and compute $\Pr(\mathcal{R}_{j})$, as defined in Equation 1. It then samples a proportional number of experiences from each shard based on $\Pr(e_{i,j}\mid\mathcal{R}_{j})$, combines them, and shuffles to form a training batch. This strategy ensures batch diversity by drawing across all shards.
@@
5.3 Integrating distributed learning in RAN systems
Figure 3: An AI architecture that integrates distributed learning in a 5G system architecture, with a learning engine realized as either a standardized or a standalone support function.
@@ 5.3 Integrating distributed learning in RAN systems
The foundational design principles of RL architectures—distributing policy evaluation (i.e., data generation) across actors while centralizing training and data management—align naturally with the hierarchical structure of RAN systems. These principles support an AI architecture in which a centralized learning engine orchestrates and provides support services to AI applications deployed in distributed RAN network functions, such as gNB centralized units, gNB distributed units, and x/rApps, which are responsible for various control and prediction tasks, as illustrated in Figure 3.
@@
The architecture in Figure 3 enables distributed learning by allowing RAN NFs that host AI-driven functionalities to collect experiences from local environments and contribute to a shared dataset for training a generalized model. Analogous to domain randomization in simulations, distributing policy evaluation across RAN NFs during training increases sample diversity, mitigates data bias, and enhances robustness against uncertainties inherent to real-world RANs.
@@
This architecture features a learning engine deployment that is agnostic to specific AI applications, while the deployment of actors (or inference functions more broadly) is use-case dependent. For example, actors replacing physical layer (L1) and medium access control layer (L2) functions—such as resource scheduling or link adaptation—could be hosted in gNB-DUs. In contrast, actors replacing networking layer (L3) functions—such as load balancing or mobility handover—could reside in gNB-CUs. To support AI-native network orchestration functions—e.g., optimizing the coverage-capacity trade-off across a network area—actors may be realized as rApps. Alternatively, actors (or inference functions) may be deployed as xApps in the nea-real-time RAN intelligent controller (RT RIC) of an ORAN system, providing insights or configuration updates to RAN NFs.
@@
To fulfill its central role and remain use-case agnostic, the learning engine can be deployed at higher layers of the RAN architecture. A cloud-based implementation provides abundant and flexible compute and storage resources to deliver AI services to distributed RAN functions. The learning engine supports a variety of services, including model training, data services (e.g., data generation control, storage, and management), and lifecycle management (LCM) operations that ensure model compliance with reliability, performance, security, ethical, and standards requirements. While LCM is centrally orchestrated, certain computational tasks—such as performance monitoring or drift detection—can be delegated to RAN NFs that execute AI models locally. This approach concentrates computationally intensive tasks, such as training and large-scale data storage, where they are most needed, thereby promoting a cost-efficient and scalable integration of AI into RAN systems.
We consider two approaches to integrating a learning engine into RAN: (a) as a standalone support function, and (b) as a standardized function.
@@
Figure 3 exemplifies a standalone learning engine coexisting with an service management and orchestration (SMO) platform in the network automation layer, communicating with actors (inference functions) hosted in RAN NFs via proprietary or extended standardized interfaces. The ML-D interface (shown in green) supports data collection from distributed RAN NFs; ML-Cd (shown in black) orchestrates training data generation for specific AI applications; and ML-Cm (shown in blue) supports artifact management and model deployment. A standalone approach offers maximum design flexibility, allowing RAN NFs and r/xApps to locally process raw data into customized training samples that meet algorithmic requirements—such as state-action-reward tuples for RL—prior to transferring them to the learning engine.
@@
A standardized approach, on the other hand, relies on standardized interfaces and functional frameworks. For example, using ORAN terminology, Figure 3 illustrates a learning engine implemented as an SMO function (SMOF), relying on other SMOFs, such as a data management entity (DME), for data collection, management, and storage. This approach mandates the use of standardized protocols and formats, which may not always align with specific AI training requirements.
@@
For instance, RAN NFs may transmit raw network observations to a DME over O1/O2 interfaces (shown in red) in accordance with interface specifications. However, these raw observations may not be directly usable for training. The DME must then process the data into training samples suited to the AI application, which are subsequently transmitted—e.g., via the R1 interface (shown in green)—to the learning engine. The primary motivation for a standardized approach is to enable inter-vendor interoperability, allowing network operators to build architectures composed of components from different vendors. However, standardizing the learning engine’s functionalities may impose significant design constraints.
@@ 6 CASE STUDY: LINK ADAPTATION
LA is a core function of modern wireless communication systems that employs adaptive modulation and coding to maximize spectral efficiency. Specifically, link adaptation (LA) optimizes the modulation order and code rate of a packet transmission to match the link capacity based on the underlying radio conditions, using receiver-side channel state information (CSI) such as channel quality indicator (CQI3rd Generation Partnership Project (3GPP). The LA parameters are then encoded into a unique value, known as the modulation and coding scheme (MCS) index, and sent to the receiver to facilitate packet decoding 3rd Generation Partnership Project (3GPP).
@@ 6.1 Outer-Loop Link Adaptation
Although several variants of this approach have been proposed Buenestado et al. (2014); Durán et al. (2015); Park et al. (2015); Blanquez-Casado et al. (2016); Ohseki and Suegara (2016); Delgado et al. (2017); Nemeth et al. (2021); Ramireddy et al. (2022); Nicolini et al. (2023), factors such as user mobility, traffic burstiness, channel aging, limited receiver-side information, and heterogeneous UE populations (e.g., with varying receiver and hardware capabilities) make LA a difficult control task to model Chen et al. (2023), rendering OLLA suboptimal.
@@ 6.2 Reinforcement Learning-Based Link Adaptation
Recognizing the limitations of traditional heuristics, there is growing interest in developing alternative, proactive solutions to LA by leveraging the predictive capabilities of machine learning (ML), and RL in particular. Numerous studies have proposed LA solutions for 5G NR networks using various Q-learning variants. Notably, Leite et al. Leite et al. (2012), Bruno et al. Bruno et al. (2014), and Mota et al. Mota et al. (2019) employed tabular Q-learning; Wu et al. Wu et al. (2020), Chen et al. Chen et al. (2023), and Kela et al. Kela et al. (2022) utilized DQN; while Geiser et al. Geiser et al. (2022) proposed an actor-critic RL algorithm for optimizing MCS selection. Other work, such as Saxena et al. Saxena et al. (2021) and Karmakar et al. Karmakar et al. (2017), has explored bandit-based approaches.
While these studies demonstrate that RL can improve LA performance over rule-based methods, they introduce new challenges that limit practical viability. For example, tabular Q-learning does not scale well with the state-action space of LA, and bandit-based approaches often rely on overly simplified problem formulations that ignore retransmissions. Most notably, many studies train and evaluate RL models using simplistic simulators that assume ideal conditions and fixed environments, without considering whether the models can generalize effectively to the diverse conditions found across radio cells in a real-world, network-wide RAN deployment.
@@ 7 MODELLING LINK ADAPTATION AS AN MDP
We jointly model the downlink LA and HARQ processes in RANs as a MDP, and present a novel DQN-based design combined with the enablers described in Section 3 to demonstrate RL generalization for RAN applications.
@@ 7.1 Episode Design
Link adaptation and HARQ operate on a per-UE and per-packet transmission basis. We formulate this problem as an episodic MDP Sutton and Barto (2018) defined by $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},p,r,\gamma,\rho_{0}\right\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p(s^{\prime}\mid s,a)$ the transition dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1)$ the discount factor, and $\rho_{0}$ the initial state distribution.
An episode models the lifespan of a UE packet in the HARQ process—from its first transmission to either a successful reception or the packet being dropped upon $N$ transmission attempts, as illustrated in Figure 4. This enables us to train a single RL policy from the collective experience generated by any UEs across the network. Each step in the episode represents the duration of a packet transmission in the HARQ process, from the selection of LA parameters (i.e., the RL action) to the reception of the associated HARQ feedback, i.e., a positive acknowledgment (ACK) or a negative acknowledgment (NACK) for successful or failed transmission, respectively. For instance, the 3GPP 5G NR system, used in our evaluations, supports at most four packet retransmissions. Hence, the episode length $N$ may range from one to five steps. Each step is characterized by a state, an action, and an associated reward, as presented next.
@@ 7.2 Reward Design
$\delta_{n}=\begin{cases}1&\text{if ACK at the $n$-th transmission attempt},\\ 0&\text{if NACK at the $n$-th transmission attempt},\end{cases}$ (3)
where $n$ denotes the packet transmission attempt, $SE_{n}$ indicates the spectral efficiency achieved at attempt $n$, and $\alpha\in\mathbb{R}$ is a scalar robustness weight penalizing retransmissions.
Optimizing exclusively for spectral efficiency may lead to excessive retransmissions, consequently lowering the effective throughput, which considers the total time required to deliver a packet. Rule-based LA algorithms, such as OLLA (see Section 6), address this issue by enforcing adherence to a predefined BLER target. Conversely, the reward formulation in Equation 2 incorporates two complementary mechanisms that encourage policies not to exploit retransmissions excessively, balancing spectral efficiency against throughput. First, when a packet is successfully transmitted (ACK) at the $n$-th attempt, $SE_{n}$ represents the spectral efficiency accounting for the cumulative time-frequency resources utilized across all $n$ transmission attempts. Second, each failed packet transmission (NACK) incurs a penalty proportional to the transmission attempt number $n$, scaled by $\alpha$.
Furthermore, the reward design in Equation 2 is agnostic to the underlying radio transmission technology, scheduling policy, and system bandwidth. This key property facilitates model generalization across heterogeneous RAN environments, enabling the deployment of a single model across nodes operating at bandwidths ranging from a few to hundreds of MHz. Specifically, formulating the reward in terms of spectral efficiency makes it independent of the transport block size (TBS) (i.e., the number of information bits in a packet) and the number of resource elements ($N_{RE}$) allocated by the scheduler. Unlike the wide-ranging values of TBS and $N_{RE}$, spectral efficiency values are standardized and bandwidth-independent, as exemplified by the MCS tables defined in the 3GPP NR specifications 3rd Generation Partnership Project (3GPP). Additionally, measuring spectral efficiency on a per-spatial-layer basis enables the training of a unified model applicable to both single- and multi-layer MIMO transmissions.
@@ 7.3 Action Space Design
The action space of the LA problem consists of a discrete set of MCS indices, represented by $\mathcal{A}=\{0,\dots,M-1\}$, each corresponding to distinct data rates defined by the communication standard. For instance, in our evaluations, we utilize the MCS indices specified by the 3GPP 5G NR specifications (e.g., Tables 5.1.3.1-1 and 5.1.3.1-2 in 3rd Generation Partnership Project (3GPP), which support modulation orders up to 64QAM and 256QAM, respectively).
Given a state $s\in\mathcal{S}$, an action $a_{m}=m$ implicitly determines the modulation order, code rate, and resulting link spectral efficiency for a packet transmission. Together with the allocated time-frequency resource blocks assigned to the UE, this action enables the calculation of the TBS for initial packet transmissions according to NR specifications 3rd Generation Partnership Project (3GPP, 3GPP). The resulting TBS is subsequently used to compute the corresponding reward, as described by Equation 2. For retransmissions, however, the TBS remains unchanged from the initial packet transmission, while the modulation order and code rate may vary.

### v3

@@ 5.1 Scaling Training in Network Simulators
In this work, we implement a large-scale distributed RL algorithm, illustrated in Figure˜2. The algorithm is based on the distributed prioritized experience replay architecture proposed in Horgan:2018, which decouples inference from learning, enables parallel inference, and significantly improves data collection capabilities. The architecture comprises three key components: a single learner operating on a graphics processing unit (GPU); multiple independent actors running concurrently on central processing units; and a centralized, yet sharded, prioritized replay memory. A sharded replay memory partitions the experience replay buffer into several smaller segments, or shards, each storing and managing a portion of the collected experiences independently—often distributed across multiple threads, processes, or machines—to enhance scalability and throughput. Each actor concurrently interacts with a group of simulations hosted by different CPUs, further accelerating the data collection process. Furthermore, each actor employs a distinct exploration policy to enrich the diversity of data. Experiences generated by each actor through interactions with its environment instances are asynchronously contributed to the replay memory shard.
@@ 5.2 Working Principles
The interaction among actors, replay memory shards, and the learner—illustrated in Figure˜2—proceeds as follows:
@@
The learner periodically queries each shard to retrieve its total priority and compute $\Pr(\mathcal{R}_{j})$, as defined in Equation˜1. It then samples a proportional number of experiences from each shard based on $\Pr(e_{i,j}\mid\mathcal{R}_{j})$, combines them, and shuffles to form a training batch. This strategy ensures batch diversity by drawing across all shards.
@@
6 Integrating Distributed Learning in RAN Systems
Figure 3: An AI architecture integrating distributed learning in a 5G RAN system, with the learning engine realized either as a standardized function within the network automation layer or as a standalone support function. A centralized learning engine performs model training, data management, and model lifecycle management, while distributed RAN network functions and r/xApps host inference (actor) functions that generate training data from local environments. Standardized and proprietary interfaces support the exchange of raw and processed data, control information, and model artifacts between the learning engine, SMO/OAM components, and RAN functions, enabling deployment across different layers of the RAN architecture.
@@ 6 Integrating Distributed Learning in RAN Systems
The foundational design principles of RL architectures—distributing policy evaluation (i.e., data generation) across actors while centralizing training and data management—align naturally with the hierarchical structure of RAN systems. These principles support an AI architecture in which a centralized learning engine orchestrates and provides support services to AI applications deployed in distributed RAN network functions, such as gNB centralized units, gNB distributed units, for various control and prediction tasks, or alternatively deployed as x/rApps in support of RAN NFs, as illustrated in Figure˜3.
@@
The architecture in Figure˜3 enables distributed learning by allowing RAN NFs that host AI-driven functionalities to collect experiences from local environments and contribute to a shared dataset for training a generalized model. Analogous to domain randomization in simulations, distributing policy evaluation across RAN NFs during training increases sample diversity, mitigates data bias, and enhances robustness against uncertainties inherent to real-world RANs. Most notably, training models directly on field data can substantially mitigate performance discrepancies induced by the sim2real gap.
@@
This architecture features a learning engine deployment that is agnostic to specific AI applications, while the deployment of actors (or inference functions more broadly) is use-case dependent. For example, actors replacing physical layer (L1) and medium access control layer (L2) functions—such as resource scheduling or link adaptation—could be hosted in gNB-DUs. In contrast, actors replacing networking layer (L3) functions—such as load balancing or mobility handover—could reside in gNB-CUs. To support AI-native network orchestration functions—e.g., optimizing the coverage-capacity trade-off across a network area—actors may be realized as rApps. Alternatively, actors (or inference functions) may be deployed as xApps in the near-real-time RAN intelligent controller (near-RT RIC) of an ORAN system, providing insights or configuration updates to RAN NFs.
@@
To fulfill its central role and remain use-case agnostic, the learning engine can be deployed at higher layers of the RAN architecture. A cloud-based implementation provides abundant and flexible compute and storage resources to deliver AI services to distributed RAN NFs underneath. The learning engine supports a variety of services, including model training, data services (e.g., data generation control, storage, and management), and model lifecycle management (LCM) operations that ensure model compliance with reliability, performance, security, ethical, and standards requirements. While LCM is centrally orchestrated, certain computational tasks—such as performance monitoring or drift detection—can be delegated to the RAN NFs that execute AI models locally. This approach concentrates computationally intensive tasks, such as training and large-scale data storage, where they are most needed, thereby promoting a cost-efficient and scalable integration of AI into RAN systems.
Figure˜3 presents two approaches to integrate a learning engine into a RAN system: (a) as a standalone support function and (b) as a standardized support function. Although Figure˜3 depicts both approaches for brevity, a practical deployment would typically adopt only one.
@@
Figure˜3 exemplifies a standalone learning engine coexisting with an service management and orchestration (SMO) platform in the network automation layer, communicating with actors (inference functions) hosted in RAN NFs via proprietary or extensions of standardized interfaces. The AI-D interface (shown in green) supports data collection from distributed RAN NFs; AI-Cd (shown in black) orchestrates training data generation for specific AI applications; and AI-Cm (shown in blue) supports artifact management and model deployment. A standalone approach offers maximum design flexibility, allowing RAN NFs and r/xApps to locally process raw data into customized training samples that meet algorithmic requirements—such as state-action-reward tuples for RL—prior to transferring them to the learning engine.
@@
A standardized approach, on the other hand, relies on standardized interfaces and functional frameworks. Using the ORAN terminology, Figure˜3 illustrates a learning engine implemented as an SMO function (SMOF), relying on other SMOFs, such as a data management entity (DME), for data collection, management, and storage. This approach mandates the use of standardized protocols and formats, which may not always align with specific AI design requirements.
@@
For instance, RAN NFs may transmit raw observations to a DME over O1/O2 interfaces (shown in red) in accordance with interface specifications. However, raw observations may not be directly usable for training. The DME must then process the data into training samples suited to the AI application, which are subsequently transmitted—e.g., via the R1 interface (shown in green)—to the learning engine. The primary motivation for a standardized approach is to enable inter-vendor interoperability, allowing network operators to build architectures composed of components from different vendors. However, standardizing the learning engine’s functionalities may pose significant design constraints.
@@ 7 Case Study: Link Adaptation
LA is a core function of modern wireless communication systems that employs adaptive modulation and coding to maximize spectral efficiency. Specifically, LA optimizes the modulation order and code rate of a packet transmission to match the link capacity based on the underlying radio conditions, using receiver-side channel state information (CSI) such as channel quality indicator (CQI) 3GPP38214. The LA parameters are then encoded into a unique value, known as the modulation and coding scheme (MCS) index, and sent to the receiver to facilitate packet decoding 3GPP38214.
@@ 7.1 Outer-Loop Link Adaptation
Although several variants of this approach have been proposed BAT+:14; DTR+:15; PDH:15; CGT+:16; OhS:16; DLM+:17; NAO:21; RGL+:22; NID:23, factors such as user mobility, traffic burstiness, channel aging, limited receiver-side information, and heterogeneous UE populations (e.g., with varying receiver and hardware capabilities) make LA a difficult control task to model Chen:23, rendering OLLA suboptimal.
@@ 7.2 Reinforcement Learning-Based Link Adaptation
Recognizing the limitations of traditional heuristics, there is growing interest in developing alternative, proactive solutions to LA by leveraging the predictive capabilities of machine learning (ML), and RL in particular. Numerous studies have proposed LA solutions for 5G NR networks using various Q-learning variants. Notably, LCV:12, BMP:14, and MAN+:19 employed tabular Q-learning; WTM:20, Chen:23, and KHV+:22 utilized DQN; while GWH+:22 proposed an actor-critic RL algorithm for optimizing MCS selection. Other work, such as saxena:21 and KCC:17, has explored bandit-based approaches.
While these studies demonstrate that RL can improve LA performance over rule-based methods, they introduce new challenges that limit practical viability. For example, tabular Q-learning does not scale well with the state-action space of LA, and bandit-based approaches often rely on overly simplified problem formulations that ignore retransmissions. Most notably, many studies train and evaluate RL models using simplistic simulators, assuming ideal conditions and fixed environments, without considering whether the models can generalize effectively to the diverse conditions found across radio cells in a real-world, network-wide RAN deployment.
@@ 8 Modelling Link Adaptation as an MDP
We jointly model the downlink LA and HARQ processes in RANs as a MDP, and present a novel DQN-based design combined with the enablers described in Section˜3 to demonstrate RL generalization for RAN applications.
@@ 8.1 Episode Design
Link adaptation and HARQ operate on a per-UE and per-packet transmission basis. We formulate this problem as an episodic MDP (Sutton1998) defined by $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},p,r,\gamma,\rho_{0}\right\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p(s^{\prime}\mid s,a)$ the transition dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1)$ the discount factor, and $\rho_{0}$ the initial state distribution.
An MDP episode models the lifespan of a UE packet in the HARQ process—from its first transmission to either a successful reception or the packet being dropped upon $N$ transmission attempts, as illustrated in Figure˜4. This enables us to train a single RL policy from the collective experience generated by any UEs across the network. Each step in the episode represents the duration of a packet transmission in the HARQ process, from the selection of LA parameters (i.e., the RL action) to the reception of the associated HARQ feedback, i.e., a positive acknowledgment (ACK) or a negative acknowledgment (NACK) for successful or failed transmission, respectively. For instance, the 3GPP 5G NR system, used in our evaluations, supports at most four packet retransmissions. Hence, the episode length $N$ may range from one to five steps. Each step is characterized by a state, an action, and an associated reward, as presented next.
@@ 8.2 Reward Design
[…]
where $n$ is a packet transmission attempt counter, $SE_{n}$ denotes the spectral efficiency achieved at attempt $n$, and $\alpha\in\mathbb{R}_{+}$ is a robustness weight penalizing retransmissions.
Optimizing exclusively for spectral efficiency may lead to excessive retransmissions, consequently lowering the effective throughput, which considers the total time required to deliver a packet. Rule-based LA algorithms, such as OLLA (see Section˜7), address this issue by enforcing adherence to a predefined BLER target. Conversely, the reward formulation in Equation˜2 incorporates two complementary mechanisms that encourage policies not to exploit retransmissions excessively, balancing spectral efficiency against throughput. First, when a packet is successfully transmitted (ACK) at the $n$-th attempt, $SE_{n}$ represents the spectral efficiency accounting for the cumulative time-frequency resources utilized across all $n$ transmission attempts. Second, each failed packet transmission (NACK) incurs a penalty proportional to the transmission attempt number $n$, scaled by $\alpha$.
Furthermore, the reward design in Equation˜2 is agnostic to the underlying radio transmission technology, scheduling policy, and system bandwidth. This key property facilitates model generalization across heterogeneous RAN environments, enabling the deployment of a single model across nodes operating at bandwidths ranging from a few to hundreds of MHz. Specifically, formulating the reward in terms of spectral efficiency makes it independent of the transport block size (TBS) (i.e., the number of information bits in a packet) and the number of resource elements ($N_{RE}$) allocated by the scheduler. Unlike the wide-ranging values of TBS and $N_{RE}$, spectral efficiency values are standardized and bandwidth-independent, as exemplified by the MCS tables defined in the 3GPP NR specifications 3GPP38214. Additionally, measuring spectral efficiency on a per-spatial-layer basis enables the training of a unified model applicable to both single- and multi-layer MIMO transmissions.
@@ 8.3 Action Space Design
The action space of the LA problem consists of a discrete set of MCS indices, represented by $\mathcal{A}=\{0,\dots,M-1\}$, each corresponding to distinct data rates defined by the communication standard. For instance, in our evaluations, we utilize the MCS indices specified by the 3GPP 5G NR specifications (e.g., Tables 5.1.3.1-1 and 5.1.3.1-2 in 3GPP38214, for modulation orders up to 64QAM and 256QAM, respectively).
Given a state $s\in\mathcal{S}$, an action $a_{m}=m$ implicitly determines the modulation order, code rate, and the targeted link spectral efficiency for a packet transmission. Together with the allocated time-frequency resource blocks assigned to the UE, this action allows us to calculate the TBS payload for the first transmission of a packet, e.g., based on NR specifications 3GPP38214; 3GPP38211. The resulting TBS is subsequently used to compute the corresponding reward, as described by Equation˜2. For retransmissions, however, the TBS remains unchanged from the initial packet transmission, while the modulation order and code rate may vary.